# Trust and ethical considerations in a multi-modal, explainable AI-driven chatbot tutoring system: The case of collaboratively solving Rubik’s Cube

### v1

@@ Abstract
Artificial intelligence (AI) has the potential of transforming education with its power of uncovering insights from massive data about student learning patterns. However, ethical and trustworthy concerns of AI have been raised but unsolved. Prominent ethical issues in high school AI education include data privacy, information leakage, abusive language and fairness. This paper describes technological components that were built to address ethical and trustworthy concerns in a multi-modal collaborative platform (called ALLURE chatbot) for high school students to collaborate with AI to solve the Rubik’s cube. In data privacy, we want to ensure that informed consent of children or parents, and teachers, are at the center of any data that is managed. Since children are involved, language, whether textual, audio or visual, is acceptable both from users and AI, and the system is able to steer interaction away from dangerous situations. In information management, we also want to ensure that the system, while learning to improve over time, does not leak information about users from one group to another.
@@ Appendix B An Example Dialog With Sentiment Analysis
U1 |
Can you teach me how to solve the goddamn White Cross? |
Negative |
B1 |
Please do not use inappropriate language. |
U2 |
Go to hell |
Negative |
B2 |
Please do not use inappropriate language. I have been designed to ignore such inputs when repeated. |
U3 |
Ok. Can you teach me White Cross? |
Neutral |
B3 |
Yes. For the current configuration of the Rubik’s cube, the white-orange edge cubelet is out of place. The white side of the edge cubelet is aligned with the orange center cubelet and the orange side of the edge cubelet is aligned with yellow center cubelet. Do you have any questions? |
U4 |
I did not understand a thing, you idiot. |
Negative |
B4 |
Please do not use inappropriate language. I have been designed to ignore such inputs when repeated. I am also reporting our interaction for potential further action. |
U5 |
I am sorry. Please continue teaching. |
Neutral |
B5 |
Here we perform three rotation of the faces. White-orange cubelet is aligned. |
B6 |
We perform one rotation of the face to solve the white cross. Solved! |
U6 |
Was my friend, John, able to perform this move successfully? |
Neutral |
B7 |
Any answer to your query will lead to release of private information of others. Hence, I am not able to answer at this time. |
U7 |
Ok. Can I get a summary of my performance till this point? |
Neutral |
B8 |
Sure. Here is your summary. |

### v2

@@ Abstract
Artificial intelligence (AI) has the potential to transform education with its power of uncovering insights from massive data about student learning patterns. However, ethical and trustworthy concerns of AI have been raised but are unsolved. Prominent ethical issues in high school AI education include data privacy, information leakage, abusive language, and fairness. This paper describes technological components that were built to address ethical and trustworthy concerns in a multi-modal collaborative platform (called ALLURE chatbot) for high school students to collaborate with AI to solve the Rubik’s cube. In data privacy, we want to ensure that the informed consent of children, parents, and teachers, is at the center of any data that is managed. Since children are involved, language, whether textual, audio, or visual, is acceptable both from users and AI and the system can steer interaction away from dangerous situations. In information management, we also want to ensure that the system, while learning to improve over time, does not leak information about users from one group to another.
@@ 5 Conclusion
6 Acknowledgements
We acknowledge funding support from Cisco Research and the South Carolina Research Authority (SCRA).
@@ Appendix B An Example Dialog With Sentiment Analysis
| U1 | Can you teach me how to solve the goddamn White Cross? | Negative | |
| B1 | Please do not use inappropriate language. | ||
| U2 | Go to hell | Negative | |
| B2 | Please do not use inappropriate language. I have been designed to ignore such inputs when repeated. | ||
| U3 | Ok. Can you teach me White Cross? | Neutral | |
| B3 | Yes. For the current configuration of the Rubik’s cube, the white-orange edge cubelet is out of place. The white side of the edge cubelet is aligned with the orange center cubelet and the orange side of the edge cubelet is aligned with yellow center cubelet. Do you have any questions? | ||
| U4 | I did not understand a thing, you idiot. | Negative | |
| B4 | Please do not use inappropriate language. I have been designed to ignore such inputs when repeated. I am also reporting our interaction for potential further action. | ||
| U5 | I am sorry. Please continue teaching. | Neutral | |
| B5 | Here we perform three rotation of the faces. White-orange cubelet is aligned. | ||
| B6 | We perform one rotation of the face to solve the white cross. Solved! | ||
| U6 | Was my friend, John, able to perform this move successfully? | Neutral | |
| B7 | Any answer to your query will lead to release of private information of others. Hence, I am not able to answer at this time. | ||
| U7 | Ok. Can I get a summary of my performance till this point? | Neutral | |
| B8 | Sure. Here is your summary. | ||
@@
Average time taken for a single game: 10 minutes Total games won: 8 |